# AmodalGen3D: Generative Amodal 3D Object Reconstruction from Sparse Unposed Views

### v1

@@ 3.1 View-Wise Cross Attention
We denote the intermediate latent as $z\in\mathbb{R}^{L\times D}$, where $L$ is the number of tokens and $D$ is the hidden dimension of the tokens. In the standard Cross Attention (denoted as $\mathcal{CA}$), the latent $z$ interacts with the encoded image feature $c$,
@@
Here, DINOv2 [oquab2023dinov2] serves as the image encoder, extracting features for input image $i$. When multiple images are available, existing approaches typically condition on one view at a time, alternating across sampling steps during the generation process. As shown in Fig. 3, this iterative conditioning disrupts consistency across views, often biasing the generation direction towards certain perspectives or causing geometric artifacts accumulation from different views, especially when sparse views differ significantly due to the 2D amodal completion.
To address the inconsistency issue, we introduce the View-Wise Cross Attention, which enables effective aggregation of 2D completion features from sparse views. Unlike the naïve sequential conditioning that alternates between views, our design allows all views to contribute simultaneously at each sampling step, promoting consistent multi-view reasoning and spatial coherence.
For each inpainted view $i^{n}$, we extract its DINO feature $c^{n}$, forming a set of encoded features $C=[c^{1},c^{2},...,c^{K}]$. We then apply Cross Attention in parallel for each view:
@@
In summary, this parallel design preserves view balance, mitigates bias toward specific perspectives, and yields more coherent latent representations across unposed sparse views during the generation process.
Figure 4: A detailed illustration of our proposed View-Wise Cross Attention and Stereo-Conditioned Cross Attention.
@@ 3.2 Stereo-Conditioned Cross Attention Module
While the View-Wise Cross Attention enhances multi-view feature fusion, it lacks explicit spatial correspondence between views. To address this, we incorporate geometric cues from a strong multi-view stereo model [wang2025pi], which provides the observed partial 3D geometry as spatial guidance in our Stereo-Conditioned Cross Attention module.
Given sparse observations $S=\{s^{1},s^{2},...,s^{K}\}$, we obtain the stereo point cloud $\mathbf{P}=\{p^{1},p^{2},...,p^{K}\}=\texttt{MVS}(S)$, where each $p^{n}$ corresponds to a point cloud reconstructed from view $s^{n}$ with pixel-level projection. We extract the partial point cloud of the target object by selecting points projected within the visibility masks:
@@
The resulting $\mathbf{P}_{O}$ is voxelized into active voxels $\mathbf{V}_{O}$ with grid size $\varphi$, then encoded by the Sparse Structure VAE, and finally transformed into geometric feature tokens using the same patchify and linear layer for transforming input noise:
@@
where $\alpha$ denotes the attention weights computed from query $Q=W_{q}z^{\prime}$ and key $K=W_{k}c_{\text{geo}}$, and $\varepsilon$ is a small constant for numerical stability in logarithmic computation. This design injects explicit geometric cues into the attention computation, improving structural consistency across unposed views and provide spatial guidance for the unobserved part.
Single-View Alternative. When only one input image is available, the partial geometry guidance can instead be derived from monocular depth estimation models [wang2025moge], providing approximate spatial context for single-view amodal reconstruction.
@@ 3.3 Data Engine & Training/Inference Proxy
Data Engine. We leverage object-centric 3D datasets from [xiang2025structured] and convert them into watertight meshes using Manifold [huang2018robust]. To simulate realistic occlusions, we generate 3D-consistent occlusion masks by randomly selecting a seed face on each mesh and iteratively expanding to its neighboring faces. This diffusion process continues outward until reaching a randomly sampled coverage ratio, producing spatially coherent and natural occlusion patterns. We then render object views and their corresponding occlusion masks, discarding camera poses with extreme angles or insufficient occlusion. Finally, completed object views are synthesized by treating the occlusion masks as inpainting regions and applying controlled mask distortions for greater diversity (see supplementary for details).
@@
We optimize the model using the standard conditional flow matching (CFM) [lipman2023flowmatchinggenerativemodeling] objective:
@@
Inference Proxy. During inference, our method can process both scene-level and object-level images. For scene-level images, we first generate the stereo point clouds for the entire scene with $S$, and then extract the partial point clouds for each object using the corresponding visibility masks $M_{vis}$. This approach reduces computational overhead when generating multiple objects within a single scene. All other procedures remain consistent with the training process. Fig. 2 shows an example of scene-level inference.
@@ 4 Experiments
Figure 5: Amodal 3D object reconstruction results on the GSO dataset. We give examples generated with 1, 2, and 4 views, respectively. Note that for FreeSplatter [xu2025freesplatter], we adopt the Hunyuan-MV  [yang2024tencent] to generate multiple views under the input of single image (first two rows).
@@ 4.2 Main Results
Comparison Methods We compare our method with various 3D object generation methods equipped with the latest 2D completion methods. First, to validate the 3D amodal completion ability with unposed sparse views of AmodalGen3D, we employ TRELLIS [xiang2025structured], Amodal3R [wu2025amodal3r], and FreeSplatter [xu2025freesplatter] for comparison. We utilize [ao2025open] to perform 2D amodal completion at sparse views. Second, to demonstrate that AmodalGen3D is also effective and can generate high-quality 3D assets in real-world scenarios, we show in-the-wild and in-the-scene qualitative results, compared with Amodal3R [wu2025amodal3r] and SceneGen [meng2025scenegen], respectively.
@@
Figure 8: In-the-wild 3D amodal reconstruction results. Single-view examples are from the COCO [lin2015microsoftcococommonobjects] dataset; sparse-view examples are from the Mip-NeRF 360 [barron2022mipnerf360] (cropped for better visual layout).
Quantitative Results. We present quantitative results tested on the GSO dataset. We first render 3D consistent masks and then generate all 1030 objects one by one. For a fair comparison, we select those views whose visible parts occupy from 30% to 70% of the whole. To validate the effectiveness and generalizability of our AmodalGen3D, we compare our method with others on different numbers of sparse view settings, i.e., 1, 2, and 4 views. As shown in Tab. LABEL:tab:_exp_metric, we employ Fréchet Inception Distance (FID) [heusel2017gans], Kernel Inception Distance (KID) [binkowski2018demystifying] to evaluate the quality of rendered images; CLIP-Score [radford2021learning] to measure the feature similarity; Minimum Matching Distance (MMD) and Coverage (COV) [nichol2022point] to evaluate the geometry quality.
@@
Qualitative Results. We here present qualitative results on the GSO dataset with simulated occlusions for amodal 3D object reconstruction. As shown in Fig. 5, our model outperforms other 3D generation methods with better geometry consistency and coherence. Moreover, with the strong 2D prior and our proposed design, AmodalGen3D shows better potential in uncommon and complex amodal object completion. We showcase more visualization results in Fig. 6, together with the 2D amodal completion results as a better demonstration of AmodalGen3D’s ability in fixing image-level structural inconsistencies in the 3D space.
View Num. FID $\downarrow$ CLIP-Score(%) $\uparrow$ MMD(‰) $\downarrow$
@@
Table 2: Ablation study on AmodalGen3D’s ability with number of input views, we test up to 20 views. The red, green indicator represents performance gain or loss compared to previous one.
@@
4.3 Ablation Study
Structure Design. We conduct ablation studies to assess the effectiveness of different module designs. Tab. 1 reports quantitative results, and Fig. 7 shows visual comparisons. With View-Wise Cross Attention, our method performs notably better under sparse views, demonstrating better rendering quality with more reasonable sparse structure as prior for texture generation. Moreover, incorporating Stereo-Conditioned Cross Attention for explicit geometry guidance yields a great benefit in the geometric metrics (MMD). As shown in Fig. 7, MVS conditioning helps AmodalGen3D produce more plausible and structurally consistent geometry, even when the 2D completions are inconsistent or contain artifacts at image level.
Number of Views. We investigate the impact of the number of input views on our method. As shown in Tab. 2, the FID, CLIP-Score, and MMD metrics remain stable or exhibit slight improvements, indicating that our method can effectively handle dense settings with up to 20 input images, which demonstrates the adaptability of our work.
2D Amodal Completion. We conduct experiments on how our model performs with different 2D amodal completion methods. As shown in Tab. 3, we can see that for different methods, our AmodalGen3D is able to produce consistent and robust results that are comparable in numerical metrics. Additionally, it is intuitive that stronger completion methods produce better results, which shows the potential and adaptability of our AmodalGen3D with more advanced techniques in the future.
Errors in MVS. In recent MVS models, the reconstructed point clouds occasionally exhibit misaligned or failing structures. Therefore, we analyze how such MVS errors propagate and affect our generation pipeline. As shown in Fig. 9, when the reconstructed stereo is geometrically implausible, our model can be less influenced by the stereo conditioning and perform self-correction with the 3D generative prior.
Figure 9: We demonstrate that AmodalGen3D can resist a certain degree of errors in the MVS reconstructed point clouds, e.g., misalignment and overlapping geometry from different views.
Figure 10: In-the-scene 3D amodal reconstruction results. Input images are from the Hypersim [roberts2021hypersim] dataset. Red and Green icons indicate different target objects (marked in the first row).
4.4 Reconstruction In-the-wild & In-the-Scene
We further evaluate our model in both in-the-wild and in-the-scene settings to assess its robustness and generalization ability. The in-the-wild cases correspond to real-world environments with complex interactions between objects, while the in-the-scene setting mainly involves structured indoor layouts, such as furniture arrangements or tabletop scenes. As shown in Fig. 8, our AmodalGen3D achieves realistic 3D amodal reconstruction across various in-the-wild conditions. In Fig. 10, we show in-the-scene reconstruction results compared with  [wu2025amodal3r] and  [meng2025scenegen], indicating the potential of our model in object-level scene reconstruction under the sparse-view and severe occlusion setting. These results demonstrate that our method can generalize across different domains and retain visible partial information more effectively while faithfully hallucinating unseen parts.
@@ 5 Conclusion
In this paper, we propose AmodalGen3D, a 3D object generator capable of reconstructing amodal 3D objects from arbitrary sparse and unposed observations. Using only object-level occlusion training data and simulated 2D amodal completion, AmodalGen3D can seamlessly integrate with 2D amodal completion methods and multi-view stereo models, enabling generalization across both synthetic and real-world scenarios and paving the way for high-quality 3D generation in diverse applications. While our amodal reconstruction relies on priors learned from generative models to infer missing regions, which may look plausible but not always match the ground-truth geometry, our approach still represents a meaningful step forward in amodal 3D reconstruction, demonstrating the potential of combining generative priors with sparse, real-world–level observations.
[…]
@@ A.3 Evaluation setting
Image-level Quality & Similarity We evaluate the image-level quality of the renderings using we employ Fréchet Inception Distance (FID) [heusel2017gans], Kernel Inception Distance (KID) [binkowski2018demystifying] as mentioned in the main paper. Additionally, we use CLIP-Score [radford2021learning] with ViT-B/32 model to calculate the feature-level cosine similarity between the rendered images and the input images. Specifically, for each object, we render 8 views using camera poses of a constant radius (2) and Field-of-View (40°), with azimuth angles of {0°, 45°, 90°, 135°, 180°, 225°, 270°, 315°} and a constant elevation angle of 15°. We calculate the score for each pair of corresponding views (generated and ground truth), and finally do a mean average for the metric score calculation.
Geometry Quality Moreover, we use Coverage (COV) and Minimum Matching Distance (MMD) with Chanmfer Distance [nichol2022point] to evaluate the geometry quality. We sample 4096 points from each ground truth and generated point cloud, which are obtained from the depth maps from different views using the farthest point sampling (FPS).

### v2

@@ 3.1 View-Wise Cross Attention
Existing multi-view conditioning strategies in generative models typically adopt either sequential cross-attention (conditioning on one view per sampling step) or naïve feature concatenation across views. As shown in Fig. 4, under sparse and unposed amodal settings, these strategies introduce two critical issues: (1) view dominance, where later-conditioned views overwrite earlier structural hypotheses, and (2) geometric drift, where inconsistencies in independently completed 2D views accumulate into unstable 3D structure.
To address this, we propose View-Wise Cross Attention, a parallelized multi-view conditioning mechanism that enforces simultaneous and balanced cross-view reasoning at every generation step. Since the base 3D generation model consistently produces objects (latents) in canonical poses regardless of the conditions’ camera poses, we exploit this pose-invariant property as a prior in designing our multi-view fusion strategy. We denote the intermediate latent as $z\in\mathbb{R}^{L\times D}$, where $L$ is the number of tokens and $D$ is the hidden dimension of the tokens. In the standard Cross Attention (denoted as $\mathcal{CA}$), the latent $z$ interacts with the encoded image feature $c$,
@@
Here, DINOv2 [oquab2023dinov2] serves as the image encoder, extracting features for input image $i$. When multiple images are available, for each 2D completion view $i^{n}$, we extract its DINO feature $c^{n}$, forming a set of encoded features $C=[c^{1},c^{2},...,c^{K}]$. We then apply Cross Attention in parallel for each view:
@@
By structurally decoupling per-view reasoning and explicitly rebalancing contributions using geometric visibility priors, View-Wise Cross Attention converts inconsistent 2D amodal completions into a coherent shared latent space, significantly improving cross-view stability under sparse observations.
@@ 3.2 Stereo-Conditioned Cross Attention Module
While prior generative models may incorporate geometric signals (e.g., depth maps or point clouds) as additional tokens, they typically treat geometry as passive conditioning features. This often results in weak structural enforcement, particularly when the geometry is partial, noisy, or unposed. To address this, we introduce Stereo-Conditioned Cross Attention, a geometry-aware attention mechanism that actively modulates the attention distribution using encoded stereo structure. Instead of merely concatenating geometric tokens, Stereo-Conditioned Cross Attention injects geometric priors directly into the attention logits via a geometry-guided gating function.
Given sparse observations $S=\{s^{1},s^{2},...,s^{K}\}$, we obtain the stereo point cloud $\mathbf{P}=\{p^{1},p^{2},...,p^{K}\}=\texttt{MVS}(S)$, where each $p^{n}$ corresponds to a point cloud reconstructed from view $s^{n}$, by lifting each pixel to the 3D space using the depth value. We extract the partial point cloud of the target object by selecting points projected within the visibility masks:
@@
The resulting $\mathbf{P}_{O}$ is voxelized into active voxels $\mathbf{V}_{O}$ with grid size $\varphi$, then encoded by the Sparse Structure VAE, and finally transformed into geometric feature tokens using the same patchify and linear layer for input transformation:
@@
where $\alpha$ denotes the attention weights computed from query $Q=W_{q}z^{\prime}$ and key $K=W_{k}c_{\text{geo}}$, and $\varepsilon$ is a small constant for logarithmic numerical stability.
By embedding 3D structure directly into the attention computation, Stereo-Conditioned Cross Attention establishes a tighter coupling between generative imagination and geometric plausibility, enforcing structural coherence even when large object regions remain unobserved.
@@ 3.3 Data Engine & Training/Inference Proxy
Data Engine. We leverage object-centric 3D datasets from [xiang2025structured] and convert them into watertight meshes using Manifold [huang2018robust]. To simulate realistic occlusions, we generate 3D-consistent occlusion masks by randomly selecting a seed face on each mesh and iteratively expanding to its neighboring faces. This process continues outward until reaching a randomly sampled coverage ratio, producing spatially coherent and natural occlusion patterns. We then render object views and their corresponding occlusion masks, discarding camera poses with extreme angles or insufficient occlusion. Finally, completed object views are synthesized by treating the occlusion masks as inpainting regions and applying controlled mask distortions for greater diversity (see supplementary for more details).
@@
We optimize the model using the standard conditional flow matching (CFM) [lipman2023flowmatchinggenerativemodeling] objective during training:
@@
Inference Proxy. During inference, our method can process both scene-level and object-level images. For scene-level images, we first generate the stereo point clouds for the entire scene with $S$, and then extract the partial point clouds for each object using the corresponding visibility masks $M_{vis}$. This approach reduces computational overhead when generating multiple objects within a single scene. All other procedures remain consistent with what is done in the training process. Fig. 2 shows an example of scene-level inference.
@@ 4 Experiments
Figure 5: Amodal 3D object generation results on the GSO dataset. We give examples generated with 1, 2, and 4 views, respectively. Note that for FreeSplatter [xu2025freesplatter], we adopt the Hunyuan-MV  [yang2024tencent] to generate multiple views under the input of single image.
Figure 7: We demonstrate that GENA3D can resist a certain degree of errors in the MVS reconstructed point clouds, e.g., misalignment and overlapping geometry from different views.
Table 1: Quantitative results on amodal 3D object generation, validated on the GSO dataset. We compare under the settings of 1, 2, and 4 views. The best and second-best results are bold and underlined, respectively. For sparse-view settings, we either apply 2D amodal completion to each view using  [ao2025open] (OAAC), or perform 2D amodal completion on a single view followed by multi-view diffusion generation (OAAC + MV)
[…]
TRELLIS [xiang2025structured] 1 view OAAC 49.68 3.21 76.79 6.08 33.37 21.1
FreeSplatter‡ [xu2025freesplatter] OAAC + MV 72.88 4.68 78.14 7.24 35.78 25.5
Amodal3R [wu2025amodal3r] N/A 39.05 0.44 80.06 5.52 38.64 46.3
SAM3D [chen2025sam] N/A 34.68 0.46 82.16 5.50 39.03 46.5
Ours OAAC 33.91 0.46 82.23 5.68 39.19 61.9
TRELLIS [xiang2025structured] 2 views OAAC 46.23 3.28 77.01 6.27 34.12 25.0
FreeSplatter [xu2025freesplatter] OAAC 91.32 6.77 74.24 10.44 27.74 12.4
Amodal3R [wu2025amodal3r] N/A 35.52 0.44 81.84 5.52 38.92 47.1
Ours OAAC 32.12 0.45 82.33 5.50 39.08 66.3
TRELLIS [xiang2025structured] 4 views OAAC 45.41 2.79 78.17 5.94 34.42 23.2
TRELLIS‡ [xiang2025structured] OAAC + MV 44.85 1.94 76.24 6.71 34.24 28.6
FreeSplatter [xu2025freesplatter] OAAC 94.10 4.64 75.58 10.48 32.31 8.4
Amodal3R [wu2025amodal3r] N/A 35.15 0.43 82.04 5.51 38.83 46.2
Ours OAAC 30.73 0.43 82.53 5.48 39.48 67.0
Table 2: The generation faithfulness to the observed visual content (regions within visibility masks $M_{vis}$). We evaluate with 1, 2, and 4 input views.
[…]
TRELLIS [xiang2025structured] 1 view 0.721 14.15 0.325
SAM3D [chen2025sam] 0.797 15.54 0.226
Amodal3R [wu2025amodal3r] 0.814 16.41 0.241
Ours 0.832 16.52 0.231
TRELLIS [xiang2025structured] 2 views 0.686 13.98 0.346
Amodal3R [wu2025amodal3r] 0.819 16.52 0.238
Ours 0.837 16.99 0.212
TRELLIS [xiang2025structured] 4 views 0.674 13.66 0.357
Amodal3R [wu2025amodal3r] 0.832 16.64 0.236
Ours 0.838 16.98 0.208
@@ 4.2 Main Results
Comparison Methods We compare our method with various 3D object generation methods equipped with the latest 2D completion methods. First, to validate the 3D amodal completion ability with unposed sparse views of GENA3D, we employ TRELLIS [xiang2025structured], Amodal3R [wu2025amodal3r], FreeSplatter [xu2025freesplatter], and the most recent SAM3D [chen2025sam] (implemented MV-SAM3D for sparse-view scenarios) for comparison. We utilize [ao2025open] to perform 2D amodal completion at sparse views. Second, to demonstrate that GENA3D is also effective and can generate high-quality 3D assets in real-world scenarios, we show in-the-wild and in-the-scene qualitative results, compared with Amodal3R [wu2025amodal3r] and SceneGen [meng2025scenegen], respectively.
@@
Quantitative Results. We present quantitative results tested on the GSO dataset. We first render 3D consistent masks and then follow the inference proxy (Sec. 3.3) to generate all 1030 objects in the dataset. For a fair comparison, we select those views whose visible parts occupy from 30% to 70% of the whole rendering. To validate the effectiveness and generalizability of our GENA3D, we compare our method with others on different numbers of sparse view settings, i.e., 1, 2, and 4 views. We note that SAM3D is evaluated solely under the single-view setting, consistent with its original design configuration and the official implementation. As shown in Tab. 1, we employ Fréchet Inception Distance (FID) [heusel2017gans], Kernel Inception Distance (KID) [binkowski2018demystifying] to evaluate the quality of rendered images; CLIP-Score [radford2021learning] to measure the feature similarity; Minimum Matching Distance (MMD) and Coverage (COV) [nichol2022point] to evaluate the geometry quality. Moreover, we evaluate the Perceptual Coherence Score (denoted as PCS) of the generated 3D objects using a VLM [bai2025qwen3]. See supplementary material for detailed explanations of applied metrics.
@@
Table 4: Ablation study on GENA3D’s ability with number of input views, we test up to 20 views. The red/green indicator represents performance gain/loss, respectively.
Num. of Views FID $\downarrow$ CLIP-Score(%) $\uparrow$ MMD(‰) $\downarrow$
@@
Qualitative Results. We present qualitative results on the GSO dataset with simulated occlusions for amodal 3D object generation. As shown in Fig. 5, our model outperforms other methods in rendering quality and geometry consistency. Moreover, by bridging 2D priors and 3D coherence, GENA3D shows better potential in uncommon and complex amodal object completion. We showcase more visualization results in Fig. 7, together with the 2D amodal completion results to demonstrate GENA3D’s ability in fixing image-level structural inconsistencies in the 3D space.
Generation Faithfulness. We also compute the average of visibility-masked SSIM [wang2004image], LPIPS [zhang2018unreasonable], and PSNR over observable regions at their corresponding views, focusing on the generation fidelity to the visible parts of sparse inputs in amodal scenarios. As shown in Tab. 2, compared to previous SoTA methods and baselines, our method achieve superior performance in terms of preserving observed visual details within the generated 3D objects.
4.3 Ablation Study
Structure Design. We conduct ablation studies to assess the effectiveness of different module designs. Tab. 3 reports quantitative results, and Fig. 8 shows visual comparisons. With View-Wise Cross Attention, our method performs notably better under sparse views, demonstrating better rendering quality with more reasonable sparse structure as prior for texture generation. Moreover, incorporating Stereo-Conditioned Cross Attention for explicit geometry guidance yields large benefits in the geometric metrics (MMD). As shown in Fig. 8, MVS conditioning helps GENA3D produce more faithful and structurally consistent geometry, even when the 2D completions are inconsistent or contain artifacts.
Number of Views. We investigate how the number of input views influence our method. As shown in Tab. 4, the provided metrics remain stable or exhibit slight improvements, indicating that our method can handle semi-dense settings with more input images, which demonstrates the adaptability of our work.
2D Amodal Completion. We conduct experiments on how our model performs with different 2D amodal completion methods. As shown in Tab. 5, we can see that for different methods, our GENA3D is able to produce consistent and robust results that are comparable in numerical metrics. Additionally, it is intuitive that stronger completion methods produce better results, which shows the potential and adaptability of our GENA3D with more advanced techniques in the future.
Errors in MVS. In recent MVS models, the reconstructed point clouds occasionally exhibit misaligned or failing structures. Therefore, we analyze how such MVS errors affect our generation pipeline. As shown in Fig. 7, when the reconstructed stereo is geometrically implausible, our model can be less influenced by the stereo conditioning and perform self-correction with the 3D generative prior.
Figure 9: In-the-wild 3D amodal generation results. Single-view examples are from the COCO [lin2015microsoftcococommonobjects] dataset; sparse-view examples are from the Mip-NeRF 360 [barron2022mipnerf360].
@@ 4.3 Ablation Study
Figure 10: In-the-scene 3D amodal generation results. Input images are from the Hypersim [roberts2021hypersim] dataset. Red and Green icons indicate different target objects.
4.4 Generation In-the-wild & In-the-Scene
We further evaluate our model in both in-the-wild and in-the-scene settings to assess its robustness and generalization ability. The in-the-wild cases correspond to real-world environments with complex interactions between objects, while the in-the-scene setting mainly involves structured indoor layouts, such as furniture arrangements or tabletop scenes. As shown in Fig. 9, our GENA3D achieves realistic 3D amodal generation across various in-the-wild conditions. In Fig. 10, we show in-the-scene generation results compared with  [wu2025amodal3r] and  [meng2025scenegen], indicating the potential of our model in object-level scene generation under the sparse-view and severe occlusion setting. These results demonstrate that with strong 2D priors and 3D coherence, our method can generalize across different domains, retain visible partial information, and faithfully hallucinate unseen parts.
@@ 5 Conclusion
In this paper, we propose GENA3D, a 3D object generator capable of generating amodal 3D objects from arbitrary sparse and partial occluded observations. Using only object-level occlusion training data and simulated 2D amodal completion, GENA3D can seamlessly integrate with 2D amodal completion methods and multi-view stereo models, enabling generalization across both synthetic and real-world scenarios and paving the way for high-quality 3D generation in diverse applications. While GENA3D relies on priors learned from generative models to infer missing regions, which may look plausible but not always match the ground-truth geometry, the approach still represents a meaningful step forward in 3D generation, demonstrating the potential of bridging strong 2D generative priors and 3D geometry coherence under real-world amodal scenarios.
@@ A.3 Evaluation setting
Image-level Quality & Similarity. We evaluate the image-level quality of the renderings using we employ Fréchet Inception Distance (FID) [heusel2017gans], Kernel Inception Distance (KID) [binkowski2018demystifying] as mentioned in the main paper. Additionally, we use CLIP-Score [radford2021learning] with ViT-B/32 model to calculate the feature-level cosine similarity between the rendered images and the input images. Specifically, for each object, we render 8 views using camera poses of a constant radius (2) and Field-of-View (40°), with azimuth angles of {0°, 45°, 90°, 135°, 180°, 225°, 270°, 315°} and a constant elevation angle of 15°. We calculate the score for each pair of corresponding views (generated and ground truth), and finally do a mean average for the metric score calculation.
Geometry Quality. Moreover, we use Coverage (COV) and Minimum Matching Distance (MMD) with Chanmfer Distance [nichol2022point] to evaluate the geometry quality. We sample 4096 points from each ground truth and generated point cloud, which are obtained from the depth maps from different views using the farthest point sampling (FPS).
Perceptual Coherence Score. Finally, we define a VLM [bai2025qwen3]-based evaluation metric to validate our method’s ability in generating visually and geometrically coherent 3D objects under amodal scenarios. We render the generated 3D objects at the same camera poses as what was done when computing the CLIP-Score. We then input the rendered views to the Qwen3-VL 32B model [bai2025qwen3] together with a systematic prompt, asking the model to rate the spatial and appearance coherence of the generated 3D object from 0 to 100. After collecting the scores for each 3D object, we do a mean average to calculate the final score. The systematic prompt is shown below:
You are an expert 3D visual quality evaluator. You will be given multiple images rendered from different viewpoints of a single 3D object. Your task is to evaluate two aspects: 1. Multi-View Coherence: Assess whether the object maintains consistent geometry, texture, color, and identity across all provided views. Penalize inconsistencies such as shape distortion, texture flickering, missing parts, or identity drift between viewpoints. 2. Appearance Quality: Assess the visual fidelity of the rendered object, including surface detail, texture sharpness, and absence of artifacts (e.g., blurriness, noise, or floaters). Scoring guidelines: - 0-20: Severe inconsistencies or artifacts, nearly unusable quality - 21-40: Noticeable issues that significantly degrade the result - 41-60: Moderate quality with some visible flaws - 61-80: Good quality with minor issues - 81-100: High quality, coherent across views, visually convincing Output format (strictly follow this): { "score": <int> } Do not include any text outside the output format.